How would you solve this equation?
2w^2=18w

Answer by stanbon(75887)

2w^2=18w
--------------
2w^2-18w = 0
Factor:
2w(w-9) = 0
2w = 0 or w-9 = 0
w = 0 or w = 9
=================
Cheers,
Stan H.